# FlowTok: Flowing Seamlessly Across Text and Image Tokens

### v2

@@ Abstract
Bridging different modalities lies at the heart of cross-modality generation. While conventional approaches treat the text modality as a conditioning signal that gradually guides the denoising process from Gaussian noise to the target image modality, we explore a much simpler paradigm—directly evolving between text and image modalities through flow matching. This requires projecting both modalities into a shared latent space, which poses a significant challenge due to their inherently different representations: text is highly semantic and encoded as 1D tokens, whereas images are spatially redundant and represented as 2D latent embeddings. To address this, we introduce FlowTok, a minimal framework that seamlessly flows across text and images by encoding images into a compact 1D token representation. Compared to prior methods, this design reduces the latent space size by 3.3 at an image resolution of 256, eliminating the need for complex conditioning mechanisms or noise scheduling. Moreover, FlowTok naturally extends to image-to-text generation under the same formulation. With its streamlined architecture centered around compact 1D tokens, FlowTok is highly memory-efficient, requires significantly fewer training resources, and achieves much faster sampling speeds—all while delivering performance comparable to state-of-the-art models. Code will be available at https://github.com/bytedance/1d-tokenizer.
@@ 1 Introduction
Bridging different modalities is essential for comprehending the diverse forms of data that represent our world, encompassing both understanding and generation. In multimodal understanding, extensive research has focused on designing architectures that project different modalities into a shared latent space [47, 55, 65, 38, 20, 83, 46, 37, 6, 24]. These approaches have significantly advanced cross-modal representation learning and real-world understanding by leveraging a common latent space between modalities.
@@
In contrast, multimodal generation (e.g., text-to-image generation) follows a different paradigm, primarily relying on the diffusion process [73, 74, 36, 61, 53], where the source modality (e.g., text) serves as a conditioning signal to guide the denoising process. Various conditioning mechanisms have been explored, including concatenation [11], cross-attention [69, 16], conditioning embeddings [61], and hybrid strategies [26, 40]. While effective, these approaches introduce substantial complexity, requiring intricate conditioning mechanisms and noise scheduling. This naturally raises an important question: Can we unify multimodal understanding and generation by enabling direct transitions within a shared latent space?
@@
To address this question, we revisit flow matching [50, 7, 54]—a modern generative framework that learns a direct path from noise to data, enabling faster convergence and accelerated sampling, leading to state-of-the-art multimodal generation results [26, 42]. Unlike diffusion models, flow matching is not constrained to using noise as the source distribution; instead, it only requires the source and target distributions to share the same shape. Pioneering works [7, 51, 72, 80, 94] have demonstrated its effectiveness in learning direct mappings within the same modality (e.g., image-to-image generation). Meanwhile, CrossFlow [52] extends flow matching to cross-modal learning by mapping text into a 2D latent space to match the shape of image embeddings, paving the way for new possibilities. However, while this approach simplifies the overall pipeline, it still operates on 2D latent representations. As a result, the additional computational overhead introduced by the text variational autoencoder [41] in CrossFlow makes it slower than modern text-to-image diffusion models like SD1.5 and SD2.1 [69], ultimately contradicting its original goal of efficiency.
@@
To this end, we introduce FlowTok, a minimal framework that enables seamless Flowing of Tokens across text and image—the two most prevalent modalities (Fig. 2). At the core of FlowTok, both text and images are encoded into compact 1D latent tokens within a unified space, enabling direct flow matching between them. On the text side, FlowTok employs a pre-trained text encoder [65] to extract initial 1D text embeddings. Since these embeddings typically reside in a higher-dimensional space than image latents, FlowTok introduces a lightweight text projector to map text embeddings into a low-dimensional variational latent space. On the image side, FlowTok builds on recent advancements in image tokenization [91, 40] to encode images into compact 1D latent tokens. Specifically, we enhance TA-TiTok [40] by integrating RoPE [76] and SwiGLU FFN [71], improving positional information handling and reconstruction quality. To enable direct flow matching, we align the number of image latent tokens in TA-TiTok to match the text encoder’s output sequence length ( for CLIP text encoder).
@@
By integrating these simple yet effective designs across text and image modalities, FlowTok represents both in the same 1D low-dimensional space with shape (77 tokens, each with 16 dimensions). This compact representation is 3.3 smaller than typical 2D flow matching shapes [75] of for image resolutions of 256. This alignment enables fast, direct flow matching and seamless evolution between the two modalities.
@@
Unlike standard flow matching models [57, 50, 54], FlowTok eliminates the need for intricate conditioning mechanisms, offering a fully self-attention-based generative model. This allows for direct flow across modalities without additional complexity. Unlike CrossFlow [52], which converts text into 2D embeddings, FlowTok retains the 1D structure of text embeddings, avoiding the need for flattening and transformation into 2D. This simplifies the framework while eliminating reliance on heavy parametric contrastive losses for semantic preservation.
@@
As a result, FlowTok offers a streamlined and resource-efficient training process. Its largest variant, FlowTok-H (1.1B), supports a batch size of 8K on 8 A100 GPUs without requiring gradient checkpointing or gradient accumulation. In contrast, recent text-to-image models of similar scale typically require 32 to 64 A100 GPUs to train with a batch size of only 2K [69, 16]. Moreover, FlowTok converges significantly faster, as shown in Fig. 3(a), with FlowTok-H completing training in just 26.1 8-A100 days—far less than SD 2.1 [69], which requires 1041.6 8-A100 GPU days.
@@
Inference is also highly efficient, with FlowTok achieving over 10 the throughput of Show-o [85] and CrossFlow [52], as shown in Fig. 3(b). This dramatically reduces computational costs, making text-to-image research far more accessible. To ensure full reproducibility, we train FlowTok exclusively on publicly available datasets, avoiding reliance on high-quality proprietary data. Remarkably, despite its minimalist design and reduced data requirements, FlowTok achieves state-of-the-art text-to-image performance.
@@
Beyond text-to-image generation, FlowTok seamlessly extends to image-to-text generation, maintaining strong performance under the same minimalist framework. We believe our work establishes a strong foundation for future research in generalized cross-modality generation. To support further advancements, all code will be released.
@@ 2 Related Work
Flow Matching. Flow matching [50, 7, 54] models generative processes by constructing a transport map between two distributions via an ordinary differential equation (ODE). It has recently gained traction as the foundation for state-of-the-art text-to-image and text-to-video synthesis models [26, 63, 18, 67], offering faster training and sampling compared to conventional diffusion methods [36, 61, 53]. Several works further optimize flow trajectories by minimizing curvature [44, 79, 64]. Despite its theoretical flexibility in handling arbitrary distributions, recent approaches primarily evolve noise into target distributions, often relying on complex control signal conditioning, which complicates the pipeline and overlooks the potential of directly transforming control signals into target distributions. In contrast, only a few works [54, 7, 51, 72, 80, 94] explore direct transport within the same modality (e.g., image-to-image [54, 94, 28]), leaving cross-modal transport (e.g., text-to-image) underexplored. In this work, we introduce FlowTok, a minimal yet effective framework that enables seamless flow across text and image modalities using 1D tokens. Unlike CrossFlow [52], which follows a similar paradigm but relies on 2D latent representations and incurs additional computational costs due to the text variational encoder, FlowTok operates within a unified, compact 1D token space. This design achieves a 3.3 compression rate in latent size, significantly reducing training costs and accelerating the sampling process, all while maintaining state-of-the-art performance.
@@
Text-to-Image Generation. Text-to-image generation has advanced rapidly in recent years, driven by various generative paradigms, including diffusion models [69, 70, 62, 16], flow matching models [26, 18, 92, 86], sequence models [89, 30, 27, 90, 68], and masked generative models [13, 84, 9, 40]. While early works in each category establish the foundation for their respective approaches, subsequent advancements across different model types have primarily emerged from three key areas: careful data collection and advanced image recaptioning for improved data quality [12, 40, 22], architectural and conditioning improvements for faster convergence and better text-image alignment [16, 26, 27, 18], and micro-conditioning for finer control over generated samples [62, 9, 40]. By contrast, this work introduces a minimalist framework FlowTok that directly maps text tokens to image tokens, eliminating the need for noise scheduling and complex conditioning mechanisms. This streamlined design enhances both efficiency and simplicity while maintaining competitive performance.
@@ 3 Preliminary
1D Visual Tokenization [91, 40] deviates from traditional 2D grid-based latent tokenization by adopting a compact 1D representation, eliminating the need to preserve the 2D spatial structure. This work focuses on continuous 1D visual tokens for flow matching. During tokenization, given an input image , the image is downscaled by a factor of , resulting in patches . These patches are concatenated with a set of latent tokens to form a sequence that is passed through a Vision Transformer (ViT) [23] encoder, , to generate embeddings. Only the embeddings corresponding to the latent tokens are retained, forming a compact 1D latent representation. This representation is modeled as a Gaussian distribution with KL divergence regularization, resulting in a compact 1D VAE representation, . In the de-tokenization phase, text guidance is applied by incorporating text embeddings generated by a pre-trained text encoder [65]. These text embeddings are projected through a linear layer to align with the channel dimensions of ViT decoder, resulting in , where is the number of context tokens predefined by the text encoder. The text embedding is then concatenated with the latent tokens and a set of mask tokens . The combined sequence is passed through the decoder , yielding the reconstructed image . Formally, with denoting concatenation, the tokenization and de-tokenization can be represented as:
Flow matching [50, 54] is a framework that learns a continuous transformation between a source distribution and a target distribution. The source distribution is not necessarily required to be Gaussian noise, though we use Gaussian noise as a concrete example below.
@@
Notably, while the source distribution is typically modeled as Gaussian noise in generative frameworks [26], the flow matching formulation generalizes to arbitrary source distributions, provided that the source and target distributions share the same shape. In FlowTok, we directly define a unified latent space for image and text modalities, treating them as both source and target distributions. This design enables seamless generation across different modalities.
@@ 4.1 Unifying Latent Space of Image and Text
Encoding Images into Compact Tokens. We build upon the core idea of TA-TiTok [40] with several enhancements to improve our image tokenizer. Specifically, we replace the original learnable 1D positional embedding with RoPE [76] to enhance TA-TiTok performance. Additionally, we substitute the MLP blocks in the Vision Transformer (ViT) [23] with SwiGLU FFN [71], which helps learn a more effective latent space [17, 88]. To align with the number of context tokens of the text encoder, we set the number of latent tokens in TA-TiTok accordingly ( for CLIP [65]). As a result, the encoder of TA-TiTok encodes each image into a compact 1D token sequence .
Transforming Texts into Compact Tokens. We use a pre-trained text encoder [65] to extract the initial text embedding , where denotes the number of channels. Notably, is typically much larger than the image latent size , as it carries richer semantic information. Since our goal is to directly flow the text embedding into the image latent space, we need to ensure both embeddings have the same shape. While we already align with through careful tuning of the image tokenizer, ensuring the image is encoded to match the length of the text tokens. The remaining challenge lies in aligning the number of channels (i.e., and ), which we resolve using a text projector. Since only the number of channels in needs adjustment while preserving its 1D shape, we employ a few simple Transformer blocks as the projector. To introduce variability in image generation from the same text, we model the projected text latents as a Gaussian distribution by applying KL divergence regularization .
A crucial aspect of text-to-image generation is ensuring that the generated image accurately reflects the input text description. Since reducing the channel dimensions of text embeddings via a learnable projector may result in semantic information loss, we introduce an auxiliary text alignment loss to preserve semantic consistency. Specifically, we employ a lightweight MLP to project into a new space for alignment. We then flatten and normalize both and along the channel dimension and compute a contrastive loss between them, inspired by CLIP [65]. Concretely, we calculate the scaled pairwise cosine similarities using a learnable temperature parameter , followed by a symmetric cross-entropy loss:
@@ 4.2 FlowTok: A General Framework for Seamless Flow Across Text and Image Tokens
Text-to-Image Generation. As shown in Fig. 4 (top), with both image and text mapped into the same latent space, FlowTok leverages vanilla flow matching [50] by stacking DiT blocks [61]. Notably, source modality (i.e., text) is directly treated as the source distribution for flow matching, removing the need for concatenation or cross-attention within the DiT blocks. This design choice further simplifies the overall framework and streamlines text-to-image generation. Combined with the compact 1D tokens introduced in Sec. 4.1, FlowTok achieves high memory efficiency, supporting a batch size of 8K on 8 A100 GPUs. Additionally, it enables fast sampling, running over faster than modern text-to-image diffusion models [69], significantly lowering the computational barrier for training large-scale text-to-image generative models.

### v3

@@ Abstract
Bridging different modalities lies at the heart of cross-modality generation. While conventional approaches treat the text modality as a conditioning signal that gradually guides the denoising process from Gaussian noise to the target image modality, we explore a much simpler paradigm—directly evolving between text and image modalities through flow matching. This requires projecting both modalities into a shared latent space, which poses a significant challenge due to their inherently different representations: text is highly semantic and encoded as 1D tokens, whereas images are spatially redundant and represented as 2D latent embeddings. To address this, we introduce FlowTok, a minimal framework that seamlessly flows across text and images by encoding images into a compact 1D token representation. Compared to prior methods, this design reduces the latent space size by 3.3 at an image resolution of 256, eliminating the need for complex conditioning mechanisms or noise scheduling. Moreover, FlowTok naturally extends to image-to-text generation under the same formulation. With its streamlined architecture centered around compact 1D tokens, FlowTok is highly memory-efficient, requires significantly fewer training resources, and achieves much faster sampling speeds—all while delivering performance comparable to state-of-the-art models. Code is available at https://github.com/TACJu/FlowTok.
@@ 1 Introduction
Bridging different modalities is essential for comprehending the diverse forms of data that represent our world, encompassing both understanding and generation. In multimodal understanding, extensive research has focused on designing architectures that project different modalities into a shared latent space [li2019visualbert, lu2019vilbert, radford2021learning, jia2021scaling, cho2021unifying, wang2021simvlm, li2022blip, jaegle2021perceiver, alayrac2022flamingo, driess2023palm, he2024a]. These approaches have significantly advanced cross-modal representation learning and real-world understanding by leveraging a common latent space between modalities.
@@
In contrast, multimodal generation (e.g., text-to-image generation) follows a different paradigm, primarily relying on the diffusion process [sohl2015deep, song2019generative, ho2020denoising, peebles2023scalable, liu2024alleviating], where the source modality (e.g., text) serves as a conditioning signal to guide the denoising process. Various conditioning mechanisms have been explored, including concatenation [bao2023all], cross-attention [rombach2022high, chen2024pixartalpha], conditioning embeddings [peebles2023scalable], and hybrid strategies [esser2024scaling, kim2025democratizing]. While effective, these approaches introduce substantial complexity, requiring intricate conditioning mechanisms and noise scheduling. This naturally raises an important question: Can we unify multimodal understanding and generation by enabling direct transitions within a shared latent space?
@@
To address this question, we revisit flow matching [lipmanflow, albergo2022building, liuflow]—a modern generative framework that learns a direct path from noise to data, enabling faster convergence and accelerated sampling, leading to state-of-the-art multimodal generation results [esser2024scaling, flux2024]. Unlike diffusion models, flow matching is not constrained to using noise as the source distribution; instead, it only requires the source and target distributions to share the same shape. Pioneering works [albergo2022building, liu2023i2sbimagetoimageschrodingerbridge, shi2023diffusion, tong2023simulation, zhou2023denoising] have demonstrated its effectiveness in learning direct mappings within the same modality (e.g., image-to-image generation). Meanwhile, CrossFlow [liu2024flowing] extends flow matching to cross-modal learning by mapping text into a 2D latent space to match the shape of image embeddings, paving the way for new possibilities. However, while this approach simplifies the overall pipeline, it still operates on 2D latent representations. As a result, the additional computational overhead introduced by the text variational autoencoder [kingma2013auto] in CrossFlow makes it slower than modern text-to-image diffusion models like SD1.5 and SD2.1 [rombach2022high], ultimately contradicting its original goal of efficiency.
@@
To this end, we introduce FlowTok, a minimal framework that enables seamless Flowing of Tokens across text and image—the two most prevalent modalities (Fig.˜2). At the core of FlowTok, both text and images are encoded into compact 1D latent tokens within a unified space, enabling direct flow matching between them. On the text side, FlowTok employs a pre-trained text encoder [radford2021learning] to extract initial 1D text embeddings. Since these embeddings typically reside in a higher-dimensional space than image latents, FlowTok introduces a lightweight text projector to map text embeddings into a low-dimensional variational latent space. On the image side, FlowTok builds on recent advancements in image tokenization [yu2024image, kim2025democratizing] to encode images into compact 1D latent tokens. Specifically, we enhance TA-TiTok [kim2025democratizing] by integrating RoPE [su2021roformer] and SwiGLU FFN [shazeer2020glu], improving positional information handling and reconstruction quality. To enable direct flow matching, we align the number of image latent tokens in TA-TiTok to match the text encoder’s output sequence length ( for CLIP text encoder).
@@
By integrating these simple yet effective designs across text and image modalities, FlowTok represents both in the same 1D low-dimensional space with shape (77 tokens, each with 16 dimensions). This compact representation is 3.3 smaller than typical 2D flow matching shapes [sdvae] of for image resolutions of 256. This alignment enables fast, direct flow matching and seamless evolution between the two modalities.
@@
Unlike standard flow matching models [ma2024sit, lipmanflow, liuflow], FlowTok eliminates the need for intricate conditioning mechanisms, offering a fully self-attention-based generative model. This allows for direct flow across modalities without additional complexity. Unlike CrossFlow [liu2024flowing], which converts text into 2D embeddings, FlowTok retains the 1D structure of text embeddings, avoiding the need for flattening and transformation into 2D. This simplifies the framework while eliminating reliance on heavy parametric contrastive losses for semantic preservation.
@@
As a result, FlowTok offers a streamlined and resource-efficient training process. Its largest variant, FlowTok-H (1.1B), supports a batch size of 8K on 8 A100 GPUs without requiring gradient checkpointing or gradient accumulation. In contrast, recent text-to-image models of similar scale typically require 32 to 64 A100 GPUs to train with a batch size of only 2K [rombach2022high, chen2024pixartalpha]. Moreover, FlowTok converges significantly faster, as shown in Fig. 3(a), with FlowTok-H completing training in just 26.1 8-A100 days—far less than SD 2.1 [rombach2022high], which requires 1041.6 8-A100 GPU days.
@@
Inference is also highly efficient, with FlowTok achieving over 10 the throughput of Show-o [xie2024show] and CrossFlow [liu2024flowing], as shown in Fig. 3(b). This dramatically reduces computational costs, making text-to-image research far more accessible. To ensure full reproducibility, we train FlowTok exclusively on publicly available datasets, avoiding reliance on high-quality proprietary data. Remarkably, despite its minimalist design and reduced data requirements, FlowTok achieves state-of-the-art text-to-image performance.
@@
Beyond text-to-image generation, FlowTok seamlessly extends to image-to-text generation, maintaining strong performance under the same minimalist framework. We believe our work establishes a strong foundation for future research in generalized cross-modality generation.
@@ 2 Related Work
Flow Matching. Flow matching [lipmanflow, albergo2022building, liuflow] models generative processes by constructing a transport map between two distributions via an ordinary differential equation (ODE). It has recently gained traction as the foundation for state-of-the-art text-to-image and text-to-video synthesis models [esser2024scaling, polyak2024moviegencastmedia, chen2025goku, ren2024flowar, ren2025grouping], offering faster training and sampling compared to conventional diffusion methods [ho2020denoising, peebles2023scalable, liu2024alleviating, liu2025revision]. Several works further optimize flow trajectories by minimizing curvature [lee2023minimizing, tong2023improving, pooladian2023multisample]. Despite its theoretical flexibility in handling arbitrary distributions, recent approaches primarily evolve noise into target distributions, often relying on complex control signal conditioning, which complicates the pipeline and overlooks the potential of directly transforming control signals into target distributions. In contrast, only a few works [liuflow, albergo2022building, liu2023i2sbimagetoimageschrodingerbridge, shi2023diffusion, tong2023simulation, zhou2023denoising] explore direct transport within the same modality (e.g., image-to-image [liuflow, zhou2023denoising, fischer2023boosting]), leaving cross-modal transport (e.g., text-to-image) underexplored. In this work, we introduce FlowTok, a minimal yet effective framework that enables seamless flow across text and image modalities using 1D tokens. Unlike CrossFlow [liu2024flowing], which follows a similar paradigm but relies on 2D latent representations and incurs additional computational costs due to the text variational encoder, FlowTok operates within a unified, compact 1D token space. This design achieves a 3.3 compression rate in latent size, significantly reducing training costs and accelerating the sampling process, all while maintaining state-of-the-art performance.
@@
Text-to-Image Generation. Text-to-image generation has advanced rapidly in recent years, driven by various generative paradigms, including diffusion models [rombach2022high, saharia2022photorealistic, podell2023sdxl, chen2024pixartalpha], flow matching models [esser2024scaling, chen2025goku, yu2025representation, yang20241, shin2025deeply], sequence models [yu2022scaling, gafni2022make, fan2024fluid, yu2024randomized, ren2025beyond], and masked generative models [chang2023muse, weber2024maskbit, bai2025meissonic, kim2025democratizing]. While early works in each category establish the foundation for their respective approaches, subsequent advancements across different model types have primarily emerged from three key areas: careful data collection and advanced image recaptioning for improved data quality [betker2023improving, kim2025democratizing, deng2025coconut], architectural and conditioning improvements for faster convergence and better text-image alignment [chen2024pixartalpha, esser2024scaling, fan2024fluid, chen2025goku], and micro-conditioning for finer control over generated samples [podell2023sdxl, bai2025meissonic, kim2025democratizing]. By contrast, this work introduces a minimalist framework FlowTok that directly maps text tokens to image tokens, eliminating the need for noise scheduling and complex conditioning mechanisms. This streamlined design enhances both efficiency and simplicity while maintaining competitive performance.
@@ 3 Preliminary
1D Visual Tokenization [yu2024image, kim2025democratizing] deviates from traditional 2D grid-based latent tokenization by adopting a compact 1D representation, eliminating the need to preserve the 2D spatial structure. This work focuses on continuous 1D visual tokens for flow matching. During tokenization, given an input image , the image is downscaled by a factor of , resulting in patches . These patches are concatenated with a set of latent tokens to form a sequence that is passed through a Vision Transformer (ViT) [dosovitskiy2020image] encoder, , to generate embeddings. Only the embeddings corresponding to the latent tokens are retained, forming a compact 1D latent representation. This representation is modeled as a Gaussian distribution with KL divergence regularization, resulting in a compact 1D VAE representation, . In the de-tokenization phase, text guidance is applied by incorporating text embeddings generated by a pre-trained text encoder [radford2021learning]. These text embeddings are projected through a linear layer to align with the channel dimensions of ViT decoder, resulting in , where is the number of context tokens predefined by the text encoder. The text embedding is then concatenated with the latent tokens and a set of mask tokens . The combined sequence is passed through the decoder , yielding the reconstructed image . Formally, with denoting concatenation, the tokenization and de-tokenization can be represented as:
Flow matching [lipmanflow, liuflow] is a framework that learns a continuous transformation between a source distribution and a target distribution. The source distribution is not necessarily required to be Gaussian noise, though we use Gaussian noise as a concrete example below.
@@
Notably, while the source distribution is typically modeled as Gaussian noise in generative frameworks [esser2024scaling], the flow matching formulation generalizes to arbitrary source distributions, provided that the source and target distributions share the same shape. In FlowTok, we directly define a unified latent space for image and text modalities, treating them as both source and target distributions. This design enables seamless generation across different modalities.
@@ 4.1 Unifying Latent Space of Image and Text
Encoding Images into Compact Tokens. We build upon the core idea of TA-TiTok [kim2025democratizing] with several enhancements to improve our image tokenizer. Specifically, we replace the original learnable 1D positional embedding with RoPE [su2021roformer] to enhance TA-TiTok performance. Additionally, we substitute the MLP blocks in the Vision Transformer (ViT) [dosovitskiy2020image] with SwiGLU FFN [shazeer2020glu], which helps learn a more effective latent space [chen2024vitamin, yao2025reconstruction]. To align with the number of context tokens of the text encoder, we set the number of latent tokens in TA-TiTok accordingly ( for CLIP [radford2021learning]). As a result, the encoder of TA-TiTok encodes each image into a compact 1D token sequence .
Transforming Texts into Compact Tokens. We use a pre-trained text encoder [radford2021learning] to extract the initial text embedding , where denotes the number of channels. Notably, is typically much larger than the image latent size , as it carries richer semantic information. Since our goal is to directly flow the text embedding into the image latent space, we need to ensure both embeddings have the same shape. While we already align with through careful tuning of the image tokenizer, ensuring the image is encoded to match the length of the text tokens. The remaining challenge lies in aligning the number of channels (i.e., and ), which we resolve using a text projector. Since only the number of channels in needs adjustment while preserving its 1D shape, we employ a few simple Transformer blocks as the projector. To introduce variability in image generation from the same text, we model the projected text latents as a Gaussian distribution by applying KL divergence regularization .
A crucial aspect of text-to-image generation is ensuring that the generated image accurately reflects the input text description. Since reducing the channel dimensions of text embeddings via a learnable projector may result in semantic information loss, we introduce an auxiliary text alignment loss to preserve semantic consistency. Specifically, we employ a lightweight MLP to project into a new space for alignment. We then flatten and normalize both and along the channel dimension and compute a contrastive loss between them, inspired by CLIP [radford2021learning]. Concretely, we calculate the scaled pairwise cosine similarities using a learnable temperature parameter , followed by a symmetric cross-entropy loss:
@@ 4.2 FlowTok: A General Framework for Seamless Flow Across Text and Image Tokens
Text-to-Image Generation. As shown in Fig. 4 (top), with both image and text mapped into the same latent space, FlowTok leverages vanilla flow matching [lipmanflow] by stacking DiT blocks [peebles2023scalable]. Notably, source modality (i.e., text) is directly treated as the source distribution for flow matching, removing the need for concatenation or cross-attention within the DiT blocks. This design choice further simplifies the overall framework and streamlines text-to-image generation. Combined with the compact 1D tokens introduced in Sec. 4.1, FlowTok achieves high memory efficiency, supporting a batch size of 8K on 8 A100 GPUs. Additionally, it enables fast sampling, running over faster than modern text-to-image diffusion models [rombach2022high], significantly lowering the computational barrier for training large-scale text-to-image generative models.